Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability Question

Expert replies
by seema19 » Mon Sep 26, 2011 4:14 am
A person has 3 children with at least one boy. Find the probablity of having at least 2 boys among the children.

Answer:[spoiler]3/4[/spoiler]

can somebody plz explain how to solve this question.
Join the discussion
Source: — Problem Solving |

by shankar.ashwin » Mon Sep 26, 2011 4:19 am
You know for sure that of the 3 children 1 is a boy, so we are concerned only about the remaining 2 children.

You need to get atleast 1 boy of the 2 to make a total of 2 boys (including the one we already know)

Again,

P(Atleast 1 Boy) = 1 - P(both girls)

= 1 - (1/2)(1/2)

3/4
seema19 wrote:A person has 3 children with at least one boy. Find the probablity of having at least 2 boys among the children.

Answer:[spoiler]3/4[/spoiler]

can somebody plz explain how to solve this question.
Join the discussion

by Anurag@Gurome » Mon Sep 26, 2011 4:56 am
seema19 wrote:A person has 3 children with at least one boy. Find the probability of having at least 2 boys among the children.

Answer:[spoiler]3/4[/spoiler]

can somebody plz explain how to solve this question.
Probability of having at least 2 boys = 1 - Probability that both are girls = 1 - 1/2 * 1/2 = 1 - 1/4 = [spoiler]3/4[/spoiler]
Anurag Mairal, Ph.D., MBA
GMAT Expert, Admissions and Career Guidance
Gurome, Inc.
1-800-566-4043 (USA)

Join Our Facebook Groups
GMAT with Gurome
https://www.facebook.com/groups/272466352793633/
Admissions with Gurome
https://www.facebook.com/groups/461459690536574/
Career Advising with Gurome
https://www.facebook.com/groups/360435787349781/
Join the discussion

by dwarzoha001 » Mon Sep 26, 2011 9:01 am
How I look at the question is this:

Chances of the three kids being:

GGG: 3!/3! = 1
GGB: 3!/2! = 3
GBB: 3!/2! = 3
BBB: 3!/3! = 1

The question states that there is already one boy, so you know GGG is not an option. So then I would just add the GBB and BBB scenarios over total possible and get [spoiler]4/7.[/spoiler] Please advise.
Join the discussion

by shankar.ashwin » Mon Sep 26, 2011 9:59 am
When you know for sure, that a kid is Boy, why would you give it a 1/2 probability again. Though there are 3 kids, you know for sure one is a boy.

What you did would be right if you didnt know if the first kid was a boy or a girl, in that case you could do;

(1+3)/8 = 4/8 (Not sure why the denominator is 7 in your Ans)
dwarzoha001 wrote:How I look at the question is this:

Chances of the three kids being:

GGG: 3!/3! = 1
GGB: 3!/2! = 3
GBB: 3!/2! = 3
BBB: 3!/3! = 1

The question states that there is already one boy, so you know GGG is not an option. So then I would just add the GBB and BBB scenarios over total possible and get [spoiler]4/7.[/spoiler] Please advise.
Join the discussion

by dwarzoha001 » Mon Sep 26, 2011 11:11 am
There are total of 8 possibilities for the gender and order of the children:

GGG BBB
GGB BBG
GBG BGB
GBB BGG

Only 1 of those possibilities is a boy not included. So from there I took the 4 possibilities of 2 boys or more, over the seven remaining scenarios. To me it's important to note that all the children are already born and at least one of the children is a boy as opposed to saying at least the first one is a boy. That's how I read into it at least.
Join the discussion

by mankey » Tue Sep 27, 2011 10:47 am
This is not clear. Please help.

Thanks
Mankey
Join the discussion